Hi,
I am not sure how to find formula to answer the question below: Help!
Angles of a triangle. The second angle of a triangular garden is four times as large as the first. The third angle is 45° less than the sum of the other two angles. Find the measure of each angle.

Answer by Mathpassionate(25)

Assume that:
1st Angle = A
2nd Angle = B
3rd Angle = C
The information provided is:

The second angle of a triangular garden is four times as large as the first.
So, B = 4A
The third angle is 45° less than the sum of the other two angles.
So, C = A+B - 45
There is a formula for triangles:
A+B+C = 180



Now, we replace the previous two formulas (B=4A) and C=A+B-45 into A+B+C = 180.
Then,
A+B+C=180
A+(4A)+(A+B-45)=180
A+(4A)+A+4A-45=180
5A+5A =180 + 45
10A = 225
A= 225/10
A = 22.5
B = 4A
B=4*(22.5)
B= 90
C= A+B-45
C= 22.5 + 90 -45
C= 67.5

So the answer of "Find the measure of each angle" is:
A = 22.5°
B= 90°
C= 67.5°